# Physics-Informed Neural Networks vs. Physics Models for Non-Invasive Glucose Monitoring: A Comparative Study Under Realistic Synthetic Conditions

### v1

@@ 4.2 Data Generation Framework Validation
Figure 3 illustrates the comprehensive ultra-realistic data generation framework developed for this study. The framework systematically incorporates all major sources of variation expected in practical prototype implementation.
Figure 3: Ultra-Realistic Data Generation Framework incorporating hardware limitations, environmental variations, physiological differences, and glucose dynamics.
The data generation process successfully achieves realistic glucose-NIR correlations of 0.211, representing a significant departure from idealized conditions where correlations can exceed 0.8. This correlation reduction reflects the combined impact of hardware noise, environmental variations, and physiological individual differences that characterize real-world implementation scenarios.
Figure 4 presents the characteristics of the generated synthetic dataset, demonstrating realistic distributions across all modeled parameters.
Figure 4: Characteristics of ultra-realistic synthetic dataset showing (a) glucose concentration distribution, (b) NIR intensity correlation with glucose, (c) environmental condition variations, and (d) physiological parameter distributions.
The glucose concentration distribution follows a realistic log-normal pattern with appropriate representation of normal, hypoglycemic, and hyperglycemic conditions. The NIR intensity correlation plot clearly demonstrates the challenging signal-to-noise conditions that practical implementations must address, with significant scatter around the underlying glucose-dependent trend.
@@ 4.3 Model Architecture Analysis
Figure 5: Detailed comparison of model architectures showing (a) Enhanced Beer–Lambert physics-based feature engineering, (b) Original PINN with Beer–Lambert constraints, (c) Optimized PINN with advanced features, (d) Full RTE PINN with complete light transport physics, (e) Selective RTE PINN with computational optimization, and (f) SDNN with shallow-deep architecture.
@@
The PINN architectures demonstrate increasing complexity from the Original PINN’s straightforward implementation to the Full RTE PINN’s comprehensive physics modeling. The trade-off between physical accuracy and computational efficiency is clearly evident in the parameter counts and inference times.
@@ 4.4 Feature Engineering Analysis
Figure 6 details the Enhanced Beer–Lambert feature engineering process, which proves crucial for achieving superior performance under realistic conditions.
Figure 6: Detailed Enhanced Beer–Lambert feature engineering process showing transformation from raw NIR intensities to 56 physics-informed features optimized for glucose detection under realistic conditions.
The feature engineering process systematically addresses the challenges of realistic glucose monitoring by incorporating glucose-specific wavelength differences that maximize signal-to-noise ratio, PMF-weighted features that compensate for individual physiological variations, environmental compensation terms that account for temperature and pressure effects, and interaction terms that capture nonlinear optical relationships.
Figure 7: Model-specific wavelength utilization heatmap. Each cell represents the relative emphasis a model places on a given NIR wavelength during prediction, with darker values indicating higher reliance. Despite 1050 nm and 1150 nm exhibiting the strongest intrinsic glucose correlation, effective glucose estimation often depends on contrasting these with low-glucose-reference bands like 850 nm and 940 nm. Only the Enhanced Beer–Lambert and Selective RTE PINN models learn to exploit this contrast structure—amplifying glucose-relevant bands while anchoring against water- and haemoglobin-dominated baselines. In contrast, unconstrained neural networks distribute attention uniformly across the spectrum, failing to identify glucose-specific optical structure.
@@ 4.5 Performance Metrics Analysis
Figure 8 presents comprehensive performance metrics across all models, clearly demonstrating the Enhanced Beer–Lambert model’s superiority under realistic conditions.
Figure 8: Comprehensive Model Performance Comparison. Six glucose prediction models are benchmarked across clinical accuracy (RMSE, MARD, Clarke Zone A, and ±15% agreement), computational efficiency (parameters and inference time), and hardware suitability. The Enhanced Beer–Lambert model achieves the best balance of accuracy and efficiency—delivering the lowest RMSE (13.6 mg/dL), highest clinical agreement (93.8% within ±15%), and requiring only 56 parameters with 0.01 ms inference time. In contrast, deep learning models without strong physics priors (e.g., SDNN, Optimized PINN) exhibit degraded accuracy and overfitting under realistic noise. While Full and Selective RTE PINNs embed physically complete modeling, their high complexity limits suitability for embedded use, highlighting the tradeoff between physical fidelity and deployment feasibility.
The performance analysis reveals several key insights. The Enhanced Beer–Lambert model achieves the best accuracy metrics while maintaining the lowest computational complexity, demonstrating the effectiveness of physics-based feature engineering. Neural network approaches show varying performance, with the Original PINN achieving competitive accuracy but at significantly higher computational cost. The RTE-based models, despite incorporating the most complete physics, do not achieve superior accuracy under realistic conditions, suggesting that the additional complexity may not be justified for practical implementations.
The computational efficiency analysis shows dramatic differences between approaches. The Enhanced Beer–Lambert model’s 56 parameters and 0.01 ms inference time make it highly suitable for embedded implementation, while the Full RTE PINN’s 1.34 million parameters and 15.2 ms inference time present significant deployment challenges.
@@ 5.1 Key Findings and Implications
The results of this comprehensive comparative analysis reveal several important findings that challenge conventional assumptions about computational approaches for non-invasive glucose monitoring. The most significant finding is that physics-based feature engineering, as implemented in the Enhanced Beer–Lambert model, can outperform complex neural network architectures under realistic conditions while maintaining computational efficiency suitable for embedded deployment.
The Enhanced Beer–Lambert model’s superior performance can be attributed to several factors. First, the physics-based feature engineering process explicitly incorporates domain knowledge about glucose-specific optical signatures, wavelength-dependent interference patterns, and individual physiological variations. This targeted approach proves more effective than relying on neural networks to automatically discover these relationships from noisy data. Second, the linear regression framework with appropriate regularization provides robustness against overfitting, which becomes particularly important when training data contains realistic noise levels. Third, the computational efficiency enables real-time implementation on resource-constrained embedded platforms.
The performance degradation observed in neural network approaches under realistic conditions highlights the challenges of applying complex models to noisy, limited datasets. The Optimized PINN, despite incorporating advanced architectural features such as residual connections and attention mechanisms, shows worse performance than the simpler Original PINN. This suggests that increased model complexity can lead to overfitting when the signal-to-noise ratio is low and training data is limited.
The RTE-based models demonstrate that incorporating more complete physics does not necessarily improve performance under realistic conditions. While the Full RTE PINN provides the most physically accurate description of light transport in tissue, its performance is limited by the quality of the underlying data and the challenges of estimating tissue optical properties from limited measurements. The computational cost of RTE modeling may not be justified for practical glucose monitoring applications.
@@ 5.2 Implications for Prototype Development
The physics-based feature engineering approach provides interpretability and robustness that are valuable for regulatory approval and clinical validation. Unlike black-box neural network approaches, the Enhanced Beer–Lambert model’s predictions can be traced back to specific optical principles and physiological parameters, facilitating understanding and validation of the measurement process.
The ultra-realistic data generation framework developed in this study provides a valuable tool for prototype development and validation. By accurately modeling hardware limitations, environmental variations, and physiological differences, the framework enables realistic performance assessment before expensive hardware prototyping. This approach can significantly reduce development time and cost while improving the likelihood of successful prototype implementation.
@@ 5.3 Limitations and Future Work
While this study provides valuable insights into computational approaches for glucose monitoring, several limitations should be acknowledged. The analysis is based on synthetic data, albeit ultra-realistic, and validation with real-world measurements will be necessary to confirm the findings. The glucose-NIR correlation of 0.211 achieved in the synthetic data represents our best estimate of realistic conditions, but actual prototype implementations may encounter additional challenges not captured in the simulation.
@@
Future research directions include validation of the Enhanced Beer–Lambert model with real-world data from prototype implementations, investigation of adaptive calibration techniques to account for individual physiological variations, exploration of hybrid approaches that combine physics-based feature engineering with selective neural network components, and development of uncertainty quantification methods to provide confidence intervals for glucose predictions.
@@ 6 Conclusions
This comprehensive comparative analysis of six computational approaches for non-invasive glucose monitoring under ultra-realistic synthetic conditions provides several important conclusions that advance the field and inform practical prototype development.
The Enhanced Beer–Lambert model with physics-based feature engineering achieves superior performance compared to all neural network approaches, demonstrating 13.6 mg/dL RMSE, 93.8% clinical accuracy, and 95.8% Clarke Error Grid Zone A classification while requiring only 56 parameters and 0.01 ms inference time. This represents a 61% improvement in RMSE compared to the SDNN baseline and establishes a new benchmark for computational efficiency in glucose monitoring applications.
The ultra-realistic synthetic data generation framework successfully models the challenging conditions of practical prototype implementation, achieving glucose-NIR correlations of 0.211 that reflect the combined impact of hardware limitations, environmental variations, and physiological individual differences. This framework provides a valuable tool for realistic performance assessment and prototype development.
Physics-based feature engineering proves more effective than complex neural network architectures under realistic conditions, suggesting that domain knowledge incorporation is crucial for robust performance in challenging measurement scenarios. The linear regression framework with appropriate regularization provides sufficient modeling capacity while maintaining robustness against overfitting.
The computational efficiency of the Enhanced Beer–Lambert model makes it highly suitable for embedded implementation on resource-constrained platforms, enabling practical deployment in portable glucose monitoring devices. The interpretability of the physics-based approach facilitates regulatory approval and clinical validation processes.
These findings establish new benchmarks for non-invasive glucose monitoring research and provide a clear path forward for practical prototype development. The Enhanced Beer–Lambert model represents a promising approach that balances accuracy, efficiency, and interpretability for real-world glucose monitoring applications.
Appendix A Model Calibration and Clinical Safety Diagnostics
Figure 10: Clarke Error Grid: The Enhanced Beer–Lambert model achieves 95.8% of predictions in Zone A (clinically accurate), with no points in Zones D or E (dangerous errors). Neural models exhibit broader spread, with accuracy ranging from 62–87%.
@@ Appendix A Model Calibration and Clinical Safety Diagnostics
Figure 13: Training dynamics of neural network models. Each subplot shows training and validation loss over epochs, highlighting convergence stability and overfitting risk. The Enhanced Beer–Lambert model is excluded from this plot as it does not require iterative training—its parameters are computed analytically via linear regression on physics-informed features.

### v2

@@ 4.2 Data Generation Framework Validation
Figure 3 illustrates the noise-stressed data generation framework developed for this study. The framework systematically incorporates major sources of variation expected in practical prototype implementation.
Figure 3: Noise-stressed data generation framework incorporating hardware limitations, environmental variations, physiological differences, and glucose dynamics.
The data generation process successfully achieves low glucose-NIR correlations of 0.211, representing a significant departure from idealized conditions where correlations can exceed 0.8. This correlation reduction reflects the combined impact of hardware noise, environmental variations, and physiological individual differences that characterize field implementation scenarios.
Figure 4 presents the characteristics of the generated synthetic dataset, demonstrating plausible distributions across all modeled parameters.
Figure 4: Characteristics of the noise-stressed synthetic dataset showing (a) glucose concentration distribution, (b) NIR intensity correlation with glucose, (c) environmental condition variations, and (d) physiological parameter distributions.
The glucose concentration distribution follows a plausible log-normal pattern with appropriate representation of normal, hypoglycemic, and hyperglycemic conditions. The NIR intensity correlation plot clearly demonstrates the challenging signal-to-noise conditions that practical implementations must address, with significant scatter around the underlying glucose-dependent trend.
@@ 4.3 Model Architecture Analysis
Figure 5: Detailed comparison of model architectures showing (a) Enhanced Beer–Lambert physics-based feature engineering, (b) Original PINN with Beer–Lambert constraints, (c) Optimized PINN with advanced features, (d) RTE-inspired PINN with physics-regularized light transport, (e) Selective RTE PINN with computational optimization, and (f) SDNN with shallow-deep architecture.
@@
The PINN architectures demonstrate increasing complexity from the Original PINN’s straightforward implementation to the RTE-inspired PINN’s physics-heavy modeling. The trade-off between physical accuracy and computational efficiency is clearly evident in the parameter counts and inference times.
@@ 4.4 Feature Engineering Analysis
Figure 6 details the Enhanced Beer–Lambert feature engineering process, which proves crucial for achieving superior performance under noise-stressed conditions.
Figure 6: Detailed Enhanced Beer–Lambert feature engineering process showing transformation from raw NIR intensities to 56 physics-informed features optimized for glucose detection under noise-stressed conditions.
The feature engineering process systematically addresses the challenges of noisy glucose monitoring by incorporating glucose-specific wavelength differences that maximize signal-to-noise ratio, PMF-weighted features that compensate for individual physiological variations, environmental compensation terms that account for temperature and pressure effects, and interaction terms that capture nonlinear optical relationships.
Figure 7: Model-specific wavelength utilization heatmap. Each cell represents the relative emphasis a model places on a given NIR wavelength during prediction, with darker values indicating higher reliance. Despite 1050 nm and 1150 nm exhibiting the strongest intrinsic glucose correlation, effective glucose estimation often depends on contrasting these with low-glucose-reference bands like 850 nm and 940 nm. Only the Enhanced Beer–Lambert and Selective RTE PINN models learn to exploit this contrast structure–amplifying glucose-relevant bands while anchoring against water- and haemoglobin-dominated baselines. In contrast, unconstrained neural networks distribute attention uniformly across the spectrum, failing to identify glucose-specific optical structure.
@@ 4.5 Performance Metrics Analysis
Figure 8 presents comprehensive performance metrics across all models, clearly demonstrating the Enhanced Beer–Lambert model’s superiority under noise-stressed conditions.
Figure 8: Comprehensive Model Performance Comparison. Six glucose prediction models are benchmarked across error metrics (RMSE, MARD), computational efficiency (parameters and inference time), and hardware suitability. The Enhanced Beer–Lambert model achieves the best balance of accuracy and efficiency–delivering the lowest RMSE (13.6 mg/dL) while requiring only 56 parameters and 0.01 ms inference time. In contrast, deep learning models without strong physics priors (e.g., SDNN, Optimized PINN) exhibit degraded accuracy under noise-stressed. While RTE-inspired and Selective RTE PINNs embed more complex physics, their higher complexity limits suitability for embedded use, highlighting the tradeoff between physical fidelity and deployment feasibility.
The performance analysis reveals several key insights. The Enhanced Beer–Lambert model achieves the best accuracy metrics while maintaining the lowest computational complexity, demonstrating the effectiveness of physics-based feature engineering. Neural network approaches show varying performance, with the Original PINN achieving competitive accuracy but at significantly higher computational cost. The RTE-based models, despite incorporating the most complete physics, do not achieve superior accuracy under noise-stressed conditions, suggesting that the additional complexity may not be justified for practical implementations.
The computational efficiency analysis shows dramatic differences between approaches. The Enhanced Beer–Lambert model’s 56 parameters and 0.01 ms inference time make it highly suitable for embedded implementation, while the RTE-inspired PINN’s 1.34 million parameters and 15.2 ms inference time present significant deployment challenges.
@@ 5.1 Key Findings and Implications
The results of this comparative analysis reveal several important findings about modeling under low SNR. The most significant finding is that physics-based feature engineering, as implemented in the Enhanced Beer–Lambert model, can outperform complex neural network architectures under noise-stressed conditions while maintaining computational efficiency suitable for embedded deployment.
The Enhanced Beer–Lambert model’s superior performance can be attributed to several factors. First, the physics-based feature engineering process explicitly incorporates domain knowledge about glucose-specific optical signatures, wavelength-dependent interference patterns, and individual physiological variations. This targeted approach proves more effective than relying on neural networks to automatically discover these relationships from noisy data. Second, the linear regression framework with appropriate regularization provides robustness against overfitting, which becomes particularly important when training data contains noise-stressed levels. Third, the computational efficiency enables real-time implementation on resource-constrained embedded platforms.
The performance degradation observed in neural network approaches under noise-stressed conditions highlights the challenges of applying complex models to noisy, limited datasets. The Optimized PINN, despite incorporating advanced architectural features, shows worse performance than the simpler Original PINN. This suggests that increased model complexity can lead to overfitting when the signal-to-noise ratio is low and training data is limited.
The RTE-based models demonstrate that incorporating more complete physics does not necessarily improve performance under noise-stressed conditions. While the RTE-inspired PINN provides the most physics-heavy description of light transport in this study, its performance is limited by the quality of the underlying data and the challenges of estimating tissue optical properties from limited measurements. The computational cost of RTE modeling may not be justified for practical glucose monitoring applications.
@@ 5.2 Implications for Prototype Development
The physics-based feature engineering approach provides interpretability and robustness. Unlike black-box neural network approaches, the Enhanced Beer–Lambert model’s predictions can be traced back to specific optical principles and physiological parameters, facilitating understanding and validation of the measurement process.
The noise-stressed data generation framework developed in this study provides a valuable tool for prototype development and validation. By modeling hardware limitations, environmental variations, and physiological differences, the framework enables performance assessment before expensive hardware prototyping. This approach can reduce development time and cost while improving the likelihood of successful prototype implementation.
@@ 5.3 Limitations and Future Work
While this study provides valuable insights into computational approaches for glucose monitoring, several limitations should be acknowledged. The analysis is based on synthetic data, albeit noise-stressed, and is not medical-grade; validation with field measurements will be necessary to confirm the findings. The glucose-NIR correlation of 0.211 achieved in the synthetic data represents our best estimate of low-SNR conditions, but actual prototype implementations may encounter additional challenges not captured in the simulation.
@@
Future research directions include validation of the Enhanced Beer–Lambert model with field data from prototype implementations, investigation of adaptive calibration techniques to account for individual physiological variations, exploration of hybrid approaches that combine physics-based feature engineering with selective neural network components, and development of uncertainty quantification methods to provide confidence intervals for glucose predictions.
@@ 6 Conclusions
This comparative analysis of six computational approaches for non-invasive glucose monitoring under noise-stressed synthetic conditions provides several conclusions that inform practical prototype development.
The Enhanced Beer–Lambert model with physics-based feature engineering achieves superior performance compared to all neural network approaches, demonstrating 13.6 mg/dL RMSE while requiring only 56 parameters and 0.01 ms inference time. This represents a 61% improvement in RMSE compared to the SDNN baseline and establishes a strong benchmark for computational efficiency in glucose monitoring applications.
The noise-stressed synthetic data generation framework successfully models challenging conditions of practical prototype implementation, achieving glucose-NIR correlations of 0.211 that reflect the combined impact of hardware limitations, environmental variations, and physiological individual differences. This framework provides a valuable tool for performance assessment and prototype development.
Physics-based feature engineering proves more effective than complex neural network architectures under noise-stressed conditions, suggesting that domain knowledge incorporation is crucial for robust performance in challenging measurement scenarios. The linear regression framework with appropriate regularization provides sufficient modeling capacity while maintaining robustness against overfitting.
The computational efficiency of the Enhanced Beer–Lambert model makes it suitable for embedded implementation on resource-constrained platforms. The interpretability of the physics-based approach supports transparent analysis of what drives predictions.
These findings provide a practical path forward for prototype development. The Enhanced Beer–Lambert model balances accuracy, efficiency, and interpretability under low-SNR conditions, showing that reliable estimation is possible even when the signal is weak and noise dominates.
Appendix A Model Calibration and Diagnostic Plots
Figure 10: Clarke Error Grid (evaluation): the Enhanced Beer–Lambert model shows tighter agreement than neural baselines under noise-stressed conditions.
@@ Appendix A Model Calibration and Diagnostic Plots
Figure 13: Training dynamics of neural network models. Each subplot shows training and validation loss over epochs, highlighting convergence stability and overfitting risk. The Enhanced Beer–Lambert model is excluded from this plot as it does not require iterative training–its parameters are computed analytically via linear regression on physics-informed features.